Hi. Here is my problem:
(a) A highway patrolman traveling at the speed limit is passed by a car going 5 mph faster than the speed limit. After one minute, the patrolman speeds up to 70 mph. How long after speeding up until the patrolman catches up with the speeding car. The speed limit is 55 mph.
(b) Same question, but this time the patrolman speeds up to a speed of v mph (v>60).
Thanks in advance!
Your response is appreciated.

Answer by solver91311(24713)



5 miles per hour is 1/12 mile per minute, so one minute after the speeder passes the cop he is 1/12 of a mile ahead. 10 miles per hour is 2/12 of a mile per minute so he catches up in 30 seconds.

If v is the new cop speed, then the speed differential is v - 60, and the speed differential in miles per minute is (v - 60)/60. Divide the 1/12 mile distance difference by the speed difference. For this situation it comes out to 5/(v - 60).

Note that this works for the original situation: 5/(70 - 60) = 5/10 = 1/2 minute = 30 seconds.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism